# When Pre-trained Visual Representations Fall Short: Limitations in Visuo-Motor Robot Learning

### v1

@@ Abstract
The integration of pre-trained visual representations (PVRs) into visuo-motor robot learning has emerged as a promising alternative to training visual encoders from scratch. However, PVRs face critical challenges in the context of policy learning, including temporal entanglement and an inability to generalise even in the presence of minor scene perturbations. These limitations hinder performance in tasks requiring temporal awareness and robustness to scene changes. This work identifies these shortcomings and proposes solutions to address them. First, we augment PVR features with temporal perception and a sense of task completion, effectively disentangling them in time. Second, we introduce a module that learns to selectively attend to task-relevant local features, enhancing robustness when evaluated on out-of-distribution scenes. Our experiments demonstrate significant performance improvements, particularly in PVRs trained with masking objectives, and validate the effectiveness of our enhancements in addressing PVR-specific limitations. Project page: tsagkas.github.io/pvrobo
Visuo-motor Robot Learning, Pre-trained Visual Representations, Behaviour Cloning
Figure 1: PCA of features from an expert demonstration in Bin Picking across PVRs (Top row: ViT models; Bottom row: ResNet models). Frame colours align with trajectory stages, suggesting feature entanglement during the gripper descent and ascent, and during the gripper stop phase. Our disentangling method improves success by $16.4$% on the Bin Picking task, up from $<$50%.
@@ 1 Introduction
Performing robust and accurate robotic manipulation from visual inputs necessitates informative and stable visual representations. The traditional paradigm for training visuo-motor policies has involved learning visual encoders from scratch alongside policy models (Levine et al., 2016). Recently, however, the adoption of pre-trained visual representations (PVRs), i.e., computer vision models trained on large and diverse visual datasets, has emerged as a compelling alternative, moving away from the tabula-rasa approach (Parisi et al., 2022). This shift is driven by three key factors: the state-of-the-art performance of PVRs in computer vision tasks, their impressive generalisation capabilities derived from training on vast datasets, and the absence of robust robot-specific foundation models capable of addressing challenges unique to robotics, such as handling diverse embodiments.
Despite the promising results of PVRs in downstream robotic applications, including affordance-based manipulation (Li et al., 2024), semantically precise tasks (Tsagkas et al., 2024), and language-guided approaches (Shen et al., 2023), their integration into visuo-motor policy learning for even basic pick-and-place tasks remains an open challenge. Crucially, training visual encoders from-scratch or fine-tuning them with in-domain data still leads to competitive performance compared to using raw PVR features or even adapted PVRs (Sharma et al., 2023; Hansen et al., 2023). Furthermore, no single PVR, or set of characteristics, has consistently delivered optimal performance across diverse tasks and environments (Majumdar et al., 2023; Hu et al., 2023). Notably, their generalisation capabilities remain underutilised, as small scene variations can destabilise policy models (Caron et al., 2021; Xie et al., 2024), c.f. Fig. 3. This limitation has reignited interest in training models from scratch with augmented datasets (Hansen et al., 2023), a strategy that is prohibitively expensive for many real-world applications.
@@
We identify these inherent characteristics of PVRs as key factors hindering visuo-motor policy learning and argue that these issues should be addressed at the feature level. Attempting to resolve the temporal entanglement problem within the policy network would limit the flexibility of PVRs in general policy architectures. For example, an LSTM policy network could be a good candidate for that (Hausknecht & Stone, 2015), but would prevent us from conditioning other SoTA approaches, e.g., diffusion policies (Chi et al., 2023). Similarly, augmenting the dataset for improving the policy’s robustness (Hansen et al., 2023) would be prohibitive for real-world robot applications, as it would require a great number of man-hours and the fine-tuning of PVR weights could affect the rich encoded knowledge.
@@
1. Identifying PVR limitations: We identify key characteristics of PVRs that hinder effective visuo-motor robot learning. Specifically, we show that they fail to encode the temporal cues and scene agnostic fine-scale visual features needed for precise manipulation tasks.
2. Temporal disentanglement: We enhance PVR features by incorporating temporal awareness and task-completion perception, without altering the policy model architecture, yielding a statistically significant improvement in downstream task performance.
3. Targeted visual features: We introduce a module that learns to directly attend to task-relevant visual cues while ignoring scene distractors. Our approach does not require dataset augmentation or re-training but instead more effectively utilises existing PVR features, particularly benefiting masked image modelling (MIM) trained PVRs.
@@ 2 Related Work
2.1 PVRs in Visuo-motor Policy Learning
In (Parisi et al., 2022), frozen PVRs were evaluated across simulated environments, outperforming models trained from scratch. Similarly, (Hu et al., 2023) showed that the utility of PVRs depends on the policy training paradigm, with behaviour cloning and inverse reinforcement learning yielding robust results, while reinforcement learning exhibited higher variability. Furthermore, (Silwal et al., 2024) provided evidence that simulation experiments (e.g., Metaworld (Yu et al., 2020) benchmark) are indicative of real world performance for PVR-based trained policies.
The dataset(s) used for pre-training plays a pivotal role in PVRs. While it was hypothesised that pre-training with video data featuring egocentric human-object interaction would be highly effective for learning features suitable for robot learning (due to their emphasis on object manipulation), research indicates that the diversity of images within the dataset is a more critical factor in successful robot learning (Dasari et al., 2023; Majumdar et al., 2023). Indeed, PVRs pre-trained on static datasets such as ImageNet (Ridnik et al., 2021) have demonstrated competitive performance, underscoring the importance of dataset variability over modality.
PVRs are favoured for their generalisation capabilities in vision tasks, but out-of-distribution generalisation remains challenging in policy deployment. (Xie et al., 2024) analysed the impact of various perturbations on PVR-based policy generalisation, while (Burns et al., 2024) identified correlations between generalisation performance and inherent model traits, such as ViTs’ segmentation ability. Conversely, (Hansen et al., 2023) found that learning from scratch with data augmentation can yield competitive results, while (Lin et al., 2024) found that adapters (Houlsby et al., 2019) can improve policy generalisation when training with diverse object instances. We focus on developing methods that achieve robustness to scene changes without relying on dataset augmentation, which can be prohibitively expensive in real-world robotics applications. Particularly, our goal is to understand the reasons PVR-based policies struggle to generalise well and argue that there is more to be achieved with untouched frozen PVRs and that their full potential in visuo-motor policy learning is yet to be revealed.
Figure 3: Visualisation of the 10 tasks used for evaluation. The first row illustrates representative scenes for all tasks, as seen in the frames from the expert demonstrations (in-domain). The second row shows how the scenes are modified by randomly altering the brightness, orientation and position of the light sourse. Similarly, the third row presents changes to the tabletop texture.
2.2 Time-informed Policy Training
In PVR-based visuo-motor policy learning, the incorporation of temporal information remains underexplored. Augmentation with temporal perception can happen either at feature level or during training time.
Feature Augmentation. While early fusion methods, such as stacking multiple frames before encoding (Karpathy et al., 2014), are common in training visual encoders from scratch, late fusion-processing frames individually and stacking their representations (Vaswani et al., 2017) has shown superior performance with fewer encoder parameters. Recent work (Shang et al., 2021) highlights that naive feature concatenation in latent space is insufficient; instead, approaches like FLARE (Shang et al., 2021) incorporate sequential embeddings and their differences, inspired by optical flow techniques. Nevertheless, concatenating sequential embeddings as input to policy networks has become standard in visuo-motor policy learning (Parisi et al., 2022) and state-of-the-art generative policies (Chi et al., 2023). However, a gap remains in leveraging PVR features, which are primarily designed for vision tasks, within this temporal framework.
Loss Function Augmentation. A major limitation of many PVRs is their inherit lack of temporal perception, as most are pre-trained on static 2D image datasets. Temporal perception can be added by employing loss functions that enforce temporal consistency during training (e.g., R3M (Nair et al., 2022) and VIP (Ma et al., 2023b)), when training with video data. However, there is no clear consensus on the superiority of this approach compared to alternatives like MIM (e.g., MVP (Xiao et al., 2022; Hu et al., 2023) and VC-1 (Majumdar et al., 2023)). This disparity suggests that existing temporal modelling strategies may be insufficient in isolation.
In subsequent experiments, we evaluate PVRs trained with temporal information and demonstrate that methods trained with a time-agnostic paradigm achieve comparable performance. We hypothesise that this limitation arises from a lack of task-completion perception, which we address by incorporating positional encoding-a fundamental mechanism in many machine learning approaches. This straightforward operation has been instrumental in the success of Transformers (Vaswani et al., 2017), implicit spatial representations (Mildenhall et al., 2020), and diffusion models (Ho et al., 2020).
2.3 Task-Relevant Feature Extraction
Downstream vision tasks often make use of the output features of PVRs. However, these features typically encode a broad range of scene information, much of which may be irrelevant to the specific task. To address this challenge, attentive probing (Chen et al., 2024; Danier et al., 2024; Bardes et al., 2024) has emerged as a popular evaluation technique, leveraging local tokens. This approach leverages a cross-attention layer with a trainable query token, treating the local features from PVRs as a sequence of key-value pairs. Unlike traditional evaluation methods such as linear probing, attentive probing has shown significantly different vision evaluation outcomes, particularly with PVRs trained using MIM approaches (e.g., MAE (He et al., 2022)), where features, such as the CLS token, often include irrelevant information. Similarly, in robot learning, task-relevant signals like joint angles may correspond to particular image regions, with unrelated cues acting as distractions. By prioritizing task-relevant signals, we show that attentive probing enhances task performance, especially in out-of-distribution scenarios.
3 Methodology
In this section, we present our approaches to enhance the deployment of PVRs in visuo-motor policy learning. First, we introduce the general behaviour cloning framework commonly employed in this domain (Section 3.1). Then, we describe a method to enrich features with temporal information, aimed at mitigating issues of temporal ambiguity (Section 3.2). Finally, we introduce our approach to selectively attend to task-relevant components of PVR features, improving their utility for robot learning tasks (Section 3.3).
3.1 Preliminaries
@@
Imitation Learning via Behaviour Cloning. We consider an expert policy $\pi^{\star}:\mathcal{P}\times\mathcal{O}\to\mathcal{A}$, which maps the current proprioceptive observation $p\in\mathcal{P}$ of a robot manipulator and a visual observation $o\in\mathcal{O}$ to a corresponding robot action $a\in\mathcal{A}$. This expert policy is used to generate a dataset of expert demonstrations, $D_{\text{exp}}$, consisting of $N$ trajectories $\mathcal{T}^{\text{e}}=\{(p_{t}^{i},o_{t}^{i},a_{t}^{i})_{t=0}^{T}\}_{i=1}^{N}$, where each trajectory captures the sequence of observations, and actions recorded over $T$ timesteps while solving a task.
The goal of behavioural cloning is to learn a policy $\pi_{\theta}$, parameterised by $\theta$, that closely imitates the expert policy by minimizing the discrepancy between its actions and the expert’s actions. This is formulated as a supervised learning problem, where the loss function measures this discrepancy across the dataset of demonstrations:
$\mathbb{E}_{(p_{t}^{i},o_{t}^{i},a_{t}^{i})\sim\mathcal{T}^{e}}\|a_{t}^{i}-\pi_{\theta}(f_{\text{PVR}}(o_{t}^{i}),p_{t}^{i})\|_{2}^{2},$ (1)
where $f_{\text{PVR}}$ represents a PVR that acts as a feature extraction function used to process visual observations $o_{t}^{i}$.
In robot learning, the decision-making process is commonly assumed to satisfy the Markov property. This assumption implies that the current observation $x_{t}=(p_{t},o_{t})$ encapsulates all the information necessary for predicting the subsequent state, i.e., $P(x_{t+1}|x_{t})=P(x_{t+1}|x_{t},x_{t-1},\dots,x_{0})$. Consequently, tasks are modelled as sequences of decisions, where each action depends solely on the current state, facilitating the application of behaviour cloning under this framework.
@@ 3.2 Temporal Disentanglement
We find that the assumption of Markovian decision-making in policies based on features extracted from frozen PVRs is frequently invalid. This arises because, at each timestep, the available information may be insufficient for the policy to confidently map the current observation to the appropriate action.
@@
Consider the example presented in Fig. 1, where PVR-features of the same pick-and-place trajectory are projected with PCA into 2D. Regardless of the PVR utilised, the extracted features seem to suffer from temporal entanglement. First, features extracted from the frames where the robot has stopped to pick up the box form a tight cluster, since the only change is the movement of the gripper fingers, which corresponds to a very small percentage of pixels. Second, as the gripper moves down and subsequently ascends, the primary visual change is the cube’s vertical displacement relative to the table. Consequently, the visual features extracted from the descent and ascent frames may differ only marginally, and only in dimensions affected by the small pixel region of the cube.
Fig. 1 also hints that including proprioceptive data in the policy input is not necessarily adequate to resolve these ambiguities. This is either because the high-dimensionality of visual features often dominates the lower-dimensional proprioceptive input, or because the robot may follow nearly identical trajectories while performing a task, further compounding the difficulty of disentangling visually similar states. At the same time, prior methods, where features from successive images were concatenated alone (Parisi et al., 2022) or along with their differences (Shang et al., 2021), have been effective to some extent. However, if frame-to-frame appearance differences are minimal, the resulting features can be very similar, and thus their differences predominantly contain near-zero values.
@@
Training a policy network to map $(p_{t},o_{t})$ to $a_{t}$ becomes difficult under these conditions. When multiple observations are nearly indistinguishable, the mapping violates the functional requirement that each input must map to exactly one output. To address these challenges, we propose a simple yet effective method to augment each observation with a temporal component by encoding the timestep index of each frame as a high-dimensional vector, using Eq. 3.2. This augmentation can temporally disentangle similar $(p_{t},o_{t})$ pairs, introducing a task progression signal into the robot state, which we argue can substantially enhance policy performance.
$\gamma(t)$ $=\left(\sin\left(\frac{2^{0}\pi t}{\text{s}^{0}}\right),\cos\left(\frac{2^{0}\pi t}{\text{s}^{0}}\right),\dots,\right.$
$\quad\left.\sin\left(\frac{2^{T-1}\pi t}{\text{s}^{T-1}}\right),\cos\left(\frac{2^{T-1}\pi t}{\text{s}^{T-1}}\right)\right)$ (2)
@@
Eq. 3.2 encodes the timestep $t$ into a high-dimensional vector $\gamma(t)$ using alternating sine and cosine functions at exponentially increasing frequencies $2^{k}$. The lower-frequency terms capture coarse temporal trends, while the higher-frequency terms provide finer temporal resolution, enabling the policy to distinguish between temporally similar states.
Figure 4: Comparison of Temporal Encoding (TE) against FLARE (Shang et al., 2021) and using no temporal augmentation on PVR features. Results (sorted by TE) show (a) per-task performance and (b) per-model performance. FLARE and TE bars indicate gains over no temporal information.
3.3 Attending to Policy Related Features

### v2

@@ Abstract
The integration of pre-trained visual representations (PVRs) into visuo-motor robot learning has emerged as a promising alternative to training visual encoders from scratch. However, PVRs face critical challenges in the context of policy learning, including temporal entanglement and an inability to generalise even in the presence of minor scene perturbations. These limitations hinder performance in tasks requiring temporal awareness and robustness to scene changes. This work identifies these shortcomings and proposes solutions to address them. First, we augment PVR features with temporal perception and a sense of task completion, effectively disentangling them in time. Second, we introduce a module that learns to selectively attend to task-relevant local features, enhancing robustness when evaluated on out-of-distribution scenes. We demonstrate significant performance improvements, particularly in PVRs trained with masking objectives, and validate the effectiveness of our enhancements in addressing PVR-specific limitations. Project page: tsagkas.github.io/pvrobo
Keywords: Pre-trained Visual Representations, Behaviour Cloning
@@ 1 Introduction
Performing robust and accurate robotic manipulation from visual inputs necessitates informative and stable visual representations. The traditional paradigm for training visuo-motor policies has involved learning visual encoders from scratch alongside policy models [1]. Recently, however, the adoption of pre-trained visual representations (PVRs), i.e., computer vision models trained on large and diverse visual datasets, has emerged as a compelling alternative, moving away from the tabula-rasa approach [2]. This shift has been driven by three key factors: the SoTA performance of PVRs in computer vision tasks, their impressive generalisation capabilities derived from training on vast datasets, and the absence of robust robot-specific foundation models capable of addressing challenges unique to robotics, such as handling diverse embodiments.
Despite the promising results of PVRs in downstream robotic applications, including affordance-based manipulation [3], semantically precise tasks [4], and language-guided approaches [5], their effective integration into visuo-motor policy learning for even basic pick-and-place tasks remains an open challenge. Crucially, training visual encoders from-scratch or fine-tuning them with in-domain data still leads to competitive performance compared to using raw PVR features or even adapted PVRs [6, 7]. Furthermore, no single PVR, or set of characteristics, has consistently delivered optimal performance across diverse tasks and environments [8, 9]. Notably, their generalisation capabilities remain underutilised, as small scene variations can destabilise policy models [10, 11], see Fig. A7. This limitation has reignited interest in training models from scratch with augmented datasets [7], a strategy that is prohibitively expensive for many real-world applications.
Figure 1: PCA of features from an expert demonstration in Bin Picking across PVRs (Top row: ViT models; Middle row: ResNet models). Frame colours align with trajectory stages, suggesting feature entanglement during the gripper descent and ascent, and during the gripper stop phase. Our disentangling method (Sec. 3.2) improves success by $16.4$% on the Bin Picking task, up from $<$50%.
@@
The aforementioned PVR characteristics are key factors hindering visuo-motor policy learning and we argue that these issues should be addressed at the feature level. Attempting to resolve the temporal entanglement problem within the policy network would limit the flexibility of PVRs in general policy architectures. For example, an LSTM policy network could be a good candidate for that [12], but would prevent us from conditioning other SoTA approaches, e.g., diffusion policies [13]. Similarly, augmenting the dataset for improving the policy’s robustness [7] would be prohibitive for real-world robot applications, as it would require a great number of man-hours and the fine-tuning of PVR weights could affect the rich encoded knowledge.
@@
1. Identifying PVR limitations. We identify key characteristics of PVRs that hinder effective visuo-motor robot learning. Specifically, we show that they fail to encode the temporal cues and scene agnostic fine-scale visual features needed for precise manipulation.
2. Temporal disentanglement. We enhance PVR features by incorporating temporal awareness and task-completion perception, without altering the policy model architecture, yielding a statistically significant improvement in downstream task performance. We also show that our method improves performance even in models that already incorporate temporal structure, whether in the input (via context with rotary positional embeddings) or the output (via action chunking [14]).
3. Targeted visual features. We introduce a module that learns to directly attend to task-relevant visual cues while ignoring scene distractors. Our approach does not require dataset augmentation or re-training but instead more effectively utilises existing PVR features, particularly benefiting masked image modelling (MIM) trained PVRs. Our approach consistently outperforms existing widely-used token pooling strategies (e.g., Spatial Softmax [15] or TokenLearner [16, 17]), under different scene variations and perturbations.
@@ 2 Related Work
PVRs in Visuo-motor Policy Learning. In [2], frozen PVRs were evaluated across simulated environments, outperforming models trained from scratch. Similarly, [9] showed that the utility of PVRs depends on the policy training paradigm, with behaviour cloning and inverse reinforcement learning yielding robust results, while reinforcement learning exhibited higher variability. Furthermore, [18] provided evidence that simulation experiments (e.g., MetaWorld [19]) are indicative of real world performance for PVR-based trained policies.
PVRs are favoured for their generalisation capabilities in vision tasks, but out-of-distribution generalisation remains challenging in policy deployment. [11] analysed the impact of various perturbations on PVR-based policy generalisation, while [20] identified correlations between generalisation performance and inherent model traits, such as ViTs’ segmentation ability. Conversely, [7] found that learning from scratch with data augmentation can yield competitive results, while [21] found that adapters [22] can improve policy generalisation when training with diverse object instances. We focus on developing methods that achieve robustness to scene changes without relying on dataset augmentation, which can be prohibitively expensive in real-world robotics applications.
Time-informed Policy Training. In PVR-based visuo-motor policy learning, the incorporation of temporal information remains underexplored. Augmentation with temporal perception can happen either at feature level or during training time. While early fusion methods, such as stacking multiple frames before encoding [23], are common in training visual encoders from scratch, late fusion (i.e., processing frames individually and stacking their representations [24]) has shown superior performance with fewer encoder parameters. Recent work [25] highlights that naive feature concatenation in latent space is insufficient; instead, approaches like FLARE [25] incorporate sequential embeddings and their differences, inspired by optical flow techniques. Nevertheless, concatenating sequential embeddings as input to policy networks has become standard in visuo-motor policy learning [2] and SoTA generative policies [13, 26]. However, a gap remains in leveraging PVR features, which are primarily designed for vision tasks, within this temporal framework.
A major limitation of many PVRs is their inherit lack of temporal perception, as most are pre-trained on static 2D image datasets. Temporal perception can be added by employing loss functions that enforce temporal consistency during training (e.g., R3M [27] and VIP [28]), when training with video data. However, there is no clear consensus on the superiority of this approach compared to alternatives like MIM (e.g., MVP [29, 9] and VC-1 [8]). This disparity suggests that existing temporal modelling strategies may be insufficient in isolation. In later experiments, we evaluate PVRs trained with temporal information and demonstrate that methods trained with a time-agnostic paradigm achieve comparable performance. We hypothesise that this limitation arises from a lack of task-completion perception, which we address by incorporating positional encoding—a fundamental mechanism in many machine learning approaches. This straightforward operation has been instrumental in the success of Transformers [30], implicit spatial representations [31], and diffusion processes [32].
Task-Relevant Feature Extraction. Downstream vision tasks often make use of the output features of PVRs. However, these features typically encode a broad range of scene information, much of which may be irrelevant to the specific task. To address this challenge, attentive probing [33, 34, 35] has emerged as a popular evaluation technique, leveraging local tokens. This approach leverages a cross-attention layer with a trainable query token, treating the local features from PVRs as a sequence of key-value pairs. Unlike traditional evaluation methods such as linear probing, attentive probing has shown significantly different vision evaluation outcomes, particularly with PVRs trained using MIM approaches (e.g., MAE [36]), where features, such as the CLS token, often include irrelevant information.
In robotics applications, pooling a sequence of observation tokens is not a novel idea and it is usually deployed for reducing the input stream length. RT-1 [17] utilised TokenLearner [16] for making the policy input more compact, thus speeding up inference time. Concurrently to our work, ICRT [37] deployed an attentive pooling module for the purpose of extracting a single state token that summarizes proprioception with visual observations. However, there is currently no consensus as to how PVR features can be effectively used such that the policy model learns to attend only to task-relevant signals, ignoring unrelated cues that act as distractors. By prioritizing task-relevant signals, we show that attentive probing enhances task performance, especially in out-of-distribution scenarios.
3 Methodology
@@
3.1 Preliminaries: Imitation Learning via Behaviour Cloning
@@
We consider an expert policy $\pi^{\star}:\mathcal{P}\times\mathcal{O}\to\mathcal{A}$, which maps a robot’s proprioceptive observation $p\in\mathcal{P}$ and visual observation $o\in\mathcal{O}$ to an action $a\in\mathcal{A}$. This policy generates a dataset $\mathcal{T}^{\text{e}}=\{(p_{t}^{i},o_{t}^{i},a_{t}^{i})_{t=0}^{T}\}_{i=1}^{N}$ of $N$ expert trajectories, where each trajectory contains $T$ steps of observations and actions for a task.
We employ behaviour cloning to learn a policy $\pi_{\theta}$, parameterised by $\theta$, to imitate $\pi^{\star}$ by minimizing the action discrepancy over demonstrations: $\mathbb{E}_{(p_{t}^{i},o_{t}^{i},a_{t}^{i})\sim\mathcal{T}^{e}}\|a_{t}^{i}-\pi_{\theta}(f_{\text{PVR}}(o_{t}^{i}),p_{t}^{i})\|_{2}^{2},$ where $f_{\text{PVR}}$ is a pre-trained visual representation (PVR) that extracts features from $o_{t}^{i}$. In visuo-motor policy learning, it is common to assume the Markov property, whereby the current observation $x_{t}=(p_{t},o_{t})$ suffices for predicting the next state: $P(x_{t+1}|x_{t})=P(x_{t+1}|x_{t},x_{t-1},\dots,x_{0})$. This allows tasks to be modelled as Markov decision processes, where each action depends only on the current state, enabling the use of behaviour cloning under this formulation.
@@ 3.2 Temporal Disentanglement
We observe that the assumption of Markovian decision-making in policies using features from frozen PVRs is often invalid. This arises because, at each timestep, the available information may be insufficient for the policy to confidently map the current observation to the appropriate action.
@@
Consider the example presented in Fig. 1, where PVR-features of the same pick-and-place trajectory are projected with PCA into 2D. Regardless of the PVR utilised, the extracted features seem to suffer from temporal entanglement. First, features extracted from the frames where the robot has stopped to pick up the box often form a tight cluster, since the only change is the movement of the gripper fingers, which corresponds to a very small percentage of pixels. Second, as the gripper moves down and subsequently ascends, the primary visual change is the cube’s vertical displacement relative to the table. Consequently, the visual features extracted from the descent and ascent frames may differ only marginally, and only in dimensions affected by the small pixel region of the cube.
@@
Training a policy network to map $(p_{t},o_{t})$ to $a_{t}$ becomes difficult under these conditions. When multiple observations are nearly indistinguishable, the mapping violates the functional requirement that each input must map to exactly one output. To address these challenges, we propose a simple yet effective method to augment each observation with a temporal component by encoding the timestep index of each frame as a high-dimensional vector, using $\gamma(t)=\Big{(}\sin\left(\frac{2^{0}\pi t}{s^{0}}\right),\cos\left(\frac{2^{0}\pi t}{s^{0}}\right),\dots,\sin\left(\frac{2^{T-1}\pi t}{s^{T-1}}\right),\cos\left(\frac{2^{T-1}\pi t}{s^{T-1}}\right)\Big{)}$ and concatenating to the policy input (see Fig. 2). This augmentation can temporally disentangle similar $(p_{t},o_{t})$ pairs, introducing a task progression signal into the robot state, which we argue can enhance policy performance.
@@
Thus, we encode each timestep $t$ into a high-dimensional vector $\gamma(t)$ using alternating sine and cosine functions at exponentially increasing frequencies $2^{k}$. The lower-frequency terms capture coarse temporal trends, while the higher-frequency terms provide finer temporal resolution, enabling the policy to distinguish between temporally similar states. Note that traditionally such embeddings encode the relative position in a transformer’s input [38], whereas here we encode the position of the embedding in the rollout.
3.3 Attending to Policy Relevant Features